OpenStudy (anonymous):

can you put this in slope intercept form x-2y+3=1 thanks

OpenStudy (anonymous):

slop intercept form is y = mx + b x - 2y + 3 = 1 -2y = -x - 3 + 1 -2y = -x - 2 (multiply the equation by -1 to make y positive) 2y = x + 2 (now divide by 2) y = (x + 2)/2 y = 1/2x + 1